# VergeIO: Depth-Aware Eye Interaction on Glasses

### v1

@@ 3.3. Hardware design
We create a smart glasses hardware prototype to ensure reliable skin contact with our chosen electrode placement (Fig. 6):
  • Electrodes. We use dry electrodes for practicality, similar to other glasses with biosignal electrodes [6, 22]. We selected SoftPulse [49] which has been used for EOG aquisition [49, 21, 50], and is biocompatible with the skin (ISO 10993-5/10).
@@
    • Ground electrode. This is part of a curved design at the end of the arm to maintains good contact with the user’s mastoid.
  • Amplifiers. We chose the BioAmp-EXG-Pill [51] for its compact analog front end (2.5 $\times$ 1 $\times$ 0.1 cm) with a 1000× signal gain, and its ability to interface with any 3.3 and 5 V MCU via ADC.
  • Microcontroller. We selected the Teensy 4.1 [52], integrated 12-bit ADC and high-performance Cortex-M7 core and enable precise, high-resolution bio-signal digitization and real-time processing.
  • Battery. We used the same 570 mAh rechargeable lithium battery as in Google Glass [53]. As discussed in Section  5.9, this battery enables continuous sensing for approximately 197 hours (8.2 days). We also integrated the Adafruit Micro Lipo Charger 4410 [54] into the glasses to enable on-board charging of the sensing system.
@@ 3.4. Motion artifact and noise removal
Motion Head tilt 30 5
@@
As EOG signals are sensitive to facial and body movements they can obscure the vergence signal [24], it is important to identify and remove these movements to avoid false positive that classify motion or noise as vergence. These movements invoke EMG (muscular) responses that can overlap in frequency with the EOG signal, resulting in signal contamination. Because of this, these signals are also not easily separable in the frequency domain.
For this reason, we leverage a data-driven, temporal segmentation approach to identify and remove segments affected by user movement. Our noise removal pipeline is designed to operate on streaming EOG data, and operate on a moving window of 2 s with an overlap of 0.1 s. We chose 2 s as it is ample time to perform a vergence movement, which typically ranges from less than 1 s [55].
@@
Signal conditioning. All signals were preprocessed using a third-order low-pass Butterworth filter with a 10 Hz cutoff, which was chosen conservatively to attenuate high-frequency noise while preserving the dominant spectral content of the EOG signal. A notch filter with 60 Hz was also adopted to suppress powerline interference.
Noise detection. To filter low-amplitude signals, particularly static noise from our data, we apply a dynamic amplitude threshold on our data and only process windows where samples exceed the threshold. Specifically, the threshold is computed by applying a median absolute deviation (MAD) filter over the first 2 seconds of each channel, which are assumed to be static, and multiplying this by a heuristic value of 4.5.
Signal classification. For each 2-second signal recording we computed statistical features for both left and right EOG channels, including signal mean, extrema, band power, wavelet energy, variance, RMS, peak-to-RMS ratio, trapezoidal integral, maximum derivative, and minimum derivative. We trained a logistic regression classifier to distinguish between vergence and non-vergence (motion/noise) segments. We evaluate the performance of our artifact removal pipeline in Sec. 5.7.
3.5. Vergence classification pipeline
@@
3:     for all $eg\in\text{Segment}(eog,2\text{s})$ s ) do
4:         if not IsArtifact($seg$then
@@
18:      $[\max-\min,\;\int g,\;(g_{e}-g_{s})/0.5,\;\overline{\Delta g},\;\mathrm{Var}(\Delta g)]$
@@
Automatic segmentation. After noise‐artifact removal, each 2 s signal window was smoothed using a Savitzky–Golay filter with a window length of 250 samples (0.5 s), which can clean the signal while preserving the features. To precisely determine the onset and offset of vergence gestures in each 2 s window, we employed the peak-detection algorithm in detecta library [56] with 30 mV as the minimum amplitude threshold, which we determined empirically from the pilot study in Sec. 3.2. Since the vergence typically range less than 1s [49], we then extracted a window centered on the detected peak, extending 0.5-seconds before and after the peak, yielding a 1 second segment.
@@
Vergence classification. Prior to model training, all features were standardized via z-score normalization to compensate for inter-user variability. The normalization also mitigates data drift due to shifts in electrode placement before and after eyeglass remounting. Using the PyCaret library [57], we automated model selection and determined that the random forest was the most effective algorithm for classifying vergence gestures.
@@ 4.1. Experimental setup
Figure 12. (a) Classification accuracy ($n=11$) for within-session (training and testing on data collected without remounting the glasses) versus cross-session (training on one session and testing on the unseen session after re-mounting glasses). P10 and P11 withdrew before session 2 due to time constraints. (b) Within-session 5-fold cross-validation accuracy per participant ($n=11$) for 2-, 3- and 4-channel configurations; P10 and P11 did not complete the 3- and 4-channel trials due to discomfort.
This study was approved by our Institutional Review Board. We recruited 11 participants (aged 23.36 ± 2.06), all were free of any ocular-related disorders, including convergence insufficiency.
All the experiments were conducted in a meeting room (Fig. 11). To trigger the six vergence gestures, we used a Brock String [58], a clinical tool commonly used to assess binocular vision and diagnose convergence insufficiency. Our string includes three colored beads positioned at the designated vergence distances: red at 30 cm, green at 70 cm, and yellow at 200 cm.
@@ 4.2. Data collection protocol
Following the standard Brock String exercise procedure [59], participants were instructed to perform each vergence gesture such that the target bead remained in sharp focus while the non-target beads appeared blurred. Our experimental protocol (Fig. 10) consisted of two sessions, each comprising ten rounds. In each round, participants performed all six vergence gestures once, cued by an audible beep, with a three-second window to complete each movement. After ten rounds, participants removed the glasses for a ten-minute break before remounting the device. For each participant, we disinfected all recording apparatus in accordance with institutional hygiene protocols.
@@ 5.1. Vergence classification
Within-Session Performance. We performed a within-session evaluation for each of the 11 participants in our study (Fig. 12a, Table 3), and achieve an average accuracy of 83.25 ± 5.48% across all participants at classifying between the 6 different vergence gestures. We perform a per-class analysis of these results as a confusion matrix (Fig. 13(a)). The matrix shows that near-to-far and far-to-near vergence movements are never confused, which is due to their distinctively different positive-going and negative-going shapes with opposing polarity. There is virtually no misclassification between the 30 $\rightarrow$ 200 cm and 70 $\rightarrow$ 200 cm transitions, or between the 200 $\rightarrow$ 70 cm and 200 $\rightarrow$ 30 cm transitions, which is partly due to the wider separation between their EOG curves. However, the 30 $\rightarrow$ 70 cm movement creates an EOG waveform that can overlap with the 30 $\rightarrow$ 200 cm and 70 $\rightarrow$ 200 cm movements, and the 70 $\rightarrow$ 30 cm movement likewise can overlap between the 200 $\rightarrow$ 70 cm and 200 $\rightarrow$ 30 cm movements.
We also consider an alternate version of our system using only four gestures which have a high degree of separability: (30 $\rightarrow$ 200 cm, 70 $\rightarrow$ 200 cm, 200 $\rightarrow$ 30 cm, and 200 $\rightarrow$ 70 cm). Our evaluation (Fig. 14a, Table 3) shows that the system achieves a within-session classification accuracy of 98.33%, and achieves near-perfect discrimination as shown in the confusion matrix in Fig. 13c.
Cross-Session Performance. When the user remounts the glasses, slight shifts in electrode placement introduce signal drift. To evaluate the robustness of our system against remounting, we trained a model on one of the user’s data collection sessions and tested it on the other unseen session. We then reversed the roles and averaged the results. Our evaluation (Fig.  12b, Table 3) shows that system performance reduces modestly from 83.25% to 79.91% after remounting. The confusion matrix in Fig. 13(b) shows increased misclassification between same-direction vergence movements. This is because remounting induces data drift. However, for the 4-gesture version of our system, cross-session performance remained stable at 98.33%, suggesting that signal drift from remounting was not substantial enough to impact classification accuracy.
@@
6-gesture accuracy 83.25 $\pm$ 5.48% 79.91 $\pm$ 7.29% 79.72 $\pm$ 3.98%
4-gesture accuracy 98.33 $\pm$ 2.04% 98.33 $\pm$ 2.20% 98.33 $\pm$ 1.67%
@@
Cross-User Performance. Here, we evaluate the performance of a pretrained model in a leave-one-user-out setup across the 11 participants in our study, where no additional user enrollment and calibration is done. Our results in Table 3 show that in this scenario, our system achieves an accuracy of 79.72% and 98.33% without any new-user enrollment for the 6-class and 4-class version of our system respectively. Additionally, we show in Fig. 14a that these accuracies surpass the personalized model trained on 3.5 minutes of new-user data, which suggests that our system can be rapidly deployed to new users without a significant decrease in system performance.
Effect of Vertical Channels. We assessed the effect of vertically placed electrodeso on system performance. Fig. 12b shows the per-participant accuracies for within-session evaluation for our proposed two‐channel configuration, a three‐ and four-channel configuration with an additional electrode pair above and below the left and right eye. 5-fold cross-validation shows that our two-channel configuration achieved 83.25% accuracy across six gestures, while the three- and four-channel configurations result in modest performance gains of 3.09% and 3.83% and accuracies of 86.57% and 87.31% respectively. These modest gains are likely because vergence is primarily, though not exclusively, a horizontal eye movement. However, all participants reported discomfort with the 3- and 4-channel configurations during data collection, two experienced ocular discomfort discontinued data collection. These findings further support our decision to exclude vertical channels from the final design.
Figure 14. (a) Effect of training duration on classification accuracy. (b) Subgroup analysis on within-session classification accuracy. (c) Performance of the motion artifact removal pipeline on EOG segments when the user is performing vergence, motion, and noise movements. (d) False positive rate under different movement scenarios: static, walking, chewing, and talking. A preamble consisting of an eyebrow raise to activate the system reduces the false positives to zero in all four conditions.
@@ 5.2. Subgroup analysis
We performed a subgroup analysis on the within-session data collected from 11 participants across 2 sessions to evaluate system performance across different demographic groups as shown in Fig. 14(b).
Sex. When comparing the classification performance by sex, female participants ($n=3$) achieved a mean accuracy of 81.67 $\pm$ 4.38%, while male participants ($n=8$) achieved 83.22 $\pm$ 6.86%. Statistical analysis using a Student’s t-test indicated that the difference was not statistically significant (t = 0.36, p = 0.72; $\alpha$ = 0.05).
Glasses. Participants who normally wore glasses were instructed to wear them with our VergeIO glasses on top of them during the study to ensure they could clearly focus on the vergence targets at 30, 70, and 200 cm. Among participants with glasses ($n=8$) and without glasses ($n=3$), the mean classification accuracies were 83.44 $\pm$ 7.11% and 81.10 $\pm$ 1.91%, respectively. No statistically significant difference was found by Student’s t-test (t = 0.54; p = 0.60; $\alpha$ = 0.05).
@@ 5.3. Baseline comparison
We perform a comparison against two baseline EOG sensing systems.
First, we compare our system with Mishra et al.’s work [5], which is focused on vergence sensing using custom aerosol-jet–printed, stretchable, skin-like biopotential electrodes using a 2, 3, and 4-channel configurations. Their 2-channel configuration differs from ours and has a horizontal channel across both eyes, a vertical channel above and below the right eye, and a ground electrode obtrusively placed on the forehead, and reports an accuracy of 34%. In contrast, our system uses an unobtrusive 2-channel configuration using standard Ag/AgCl electrodes with the ground on the mastoid and achieves an accuracy of 83.25%. While their 3-channel and 4-channel configurations have increased accuracies of 87 and 91% respectively, additional electrodes bring about discomfort which negatively affect the user experience.
Mishra et al. [5] VergeIO (ours)
Number of electrodes 5, 5, 7 4
Number of channels 2, 3, 3 2
In-session error 34%, 87%, 91% 83.25%
Cross-session error Not specified 79.91%
Cross-user error Not specified 79.72%
@@
Table 4. Comparison between our system and the baseline proposed by Mishra et al.  [5]
Second, we compare our system’s performance against the EOG electrode layout in the commercial JINS MEME glasses [6] (Fig. 5). Here, one participant (P4) wore electrodes at the union of both our system’s and JINS MEME’s electrode positions. We then collected data across ten rounds, each including all six vergence gestures. Five‐fold cross‐validation was performed independently on both setups. Our system achieves an accuracy of 88.33 $\pm$ 1.13%, while the JINS MEME achieves a lower accuracy of 50.0 $\pm$ 1.29%. This is in line with our prior SNR analysis showing that our vergence waveforms are on average 7.3 dB higher than those from JINS MEME (Fig. 7).
5.4. Classifying vergence and non-vergence eye movements
We evaluate whether our design can classify vergence, as well as other eye movements. We collect a dataset including the four vergence gestures with high separability, along with seven additional eye movements: left and right winks, blinks, and four directional gazes (up, down, left, right). We collected 14 recordings for each class, and performed 5-fold cross validation using logistic regression. Our system achieved an accuracy of 96.1 $\pm$ 2.4%.
A few misclassifications were observed between upward and downward gaze, which is expected as our design intentionally omits vertical electrode channels to avoid user discomfort. However, our system still captures some vertical eye activity as eye movements are not perfectly directional and there are slight asymmetries in electrode placement. These results show that our electrode configuration does not tradeoff accuracy at detecting standard eye movements sensed in prior EOG systems, and can detect them robustly along with vergence gestures.

### v2

@@ 3.3. Hardware design
We create a glasses hardware prototype to ensure reliable skin contact of our chosen electrode placement (Fig. 6):
  • Electrodes. We use dry electrodes for practicality, similar to other glasses with biosignal electrodes (jin, 2026; Kosmyna et al., 2019). We selected SoftPulse (Chen et al., 2014) which has been used for EOG aquisition (Chen et al., 2014; Frey et al., 2024; Heijden et al., 2024), and is biocompatible with the skin (ISO 10993-5/10).
@@
    • Ground electrode. Curved coating at the end of the arm to maintains good contact with the user’s mastoid.
  • Amplifiers. We chose the BioAmp-EXG-Pill (amp, 2026) for its compact analog front end (2.5 $\times$ 1 $\times$ 0.1 cm) with a 1000× signal gain, and its ability to interface with any 3.3 and 5 V MCU via ADC.
  • Microcontroller. We selected the Teensy 4.1 (tee, 2026), integrated 12-bit ADC and high-performance Cortex-M7 core and enable precise, high-resolution bio-signal digitization and real-time processing.
  • Battery. We used the same 570 mAh rechargeable lithium battery as in Google Glass (ggl, 2026b). As discussed in Section  5.9, this battery enables continuous sensing for approximately 197 hours (8.2 days). We also integrated the Adafruit Micro Lipo Charger 4410 (ada, 2026) into the glasses to enable on-board charging of the sensing system.
@@ 3.4. Motion artifact and noise removal
Motion Head turn 30 5
@@
Given that EOG signals are highly susceptible to facial and body movements(Rostaminia et al., 2019), the identification and removal of such artifacts are imperative to prevent false positives. Since the resulting EMG (muscular) responses overlap with the EOG signals in frequency, they cannot be easily isolated in the frequency domain.
For this reason, we leverage a temporal, data-driven segmentation approach to identify and remove segments affected by user movement. Our noise removal pipeline is designed to operate on streaming EOG data, and operate on a moving window of 2 s with an overlap of 0.1 s. We chose 2 s as it is ample time to perform a vergence movement, which typically ranges from less than 1 s (Tyler et al., 2012).
@@
Signal Conditioning. All signals were preprocessed using a third-order low-pass Butterworth filter with a 10 Hz cutoff, which was chosen to attenuate high-frequency noise while preserving the dominant spectral content of the EOG signal (Mishra et al., 2020). A notch filter working at 60 Hz was also adopted to suppress powerline interference.
Silence Removal. To filter low-amplitude signals, particularly static noise from our data, we apply a dynamic amplitude threshold on our data and only process windows where samples exceed the threshold. Specifically, the threshold is computed by applying a median absolute deviation (MAD) filter over the first 2 seconds of each channel, which are assumed to be static, and multiplying this by a heuristic value of 4.5.
Artifact Detection and Removal. For each 2-second signal recording we computed statistical features for both left and right EOG channels, including signal mean, extrema, band power, wavelet energy, variance, RMS, peak-to-RMS ratio, trapezoidal integral, maximum derivative, and minimum derivative. We trained a logistic regression classifier to distinguish between vergence and non-vergence (motion/noise) segments and remove the latter. We evaluate the performance of our artifact removal pipeline in Sec. 5.7.
@@
3:  for all $seg\in\text{Segment}(eog,2\text{s})$ do
4:   if not IsArtifact(segseg) then
@@
18:   $\max-\min,\;\int g,\;(g_{e}-g_{s})/0.5,\;\overline{\Delta g},\;\mathrm{Var}(\Delta g)]$
@@
3.5. Vergence classification pipeline
Automatic Segmentation. After noise‐artifact removal, each 2 s signal window was smoothed using a Savitzky–Golay filter (Savitzky and Golay, 1964) with a window length of 250 samples (0.5 s), which can clean the signal while preserving the features. Compared with moving averages or wavelet denoising, Savitzky–Golay filter preserves edges more faithfully and is computationally efficient for real-time use (Acharya et al., 2016; Abdullah et al., 2023). To precisely determine the onset and offset of vergence gestures in each 2 s window, we employed the peak-detection algorithm in detecta library (Duarte, 2021) with 30 mV as the minimum amplitude threshold, which we determined empirically from the pilot study in Sec. 3.2. Since the vergence typically range less than 1s (Chen et al., 2014), we then extracted a window centered on the detected peak, extending 0.5-seconds before and after the peak, yielding a 1 second segment.
@@ 3.5. Vergence classification pipeline
Vergence Classification. Prior to model training, all features were standardized via z-score normalization to compensate for inter-user variability. The normalization also mitigates data drift due to shifts in electrode placement before and after eyeglass remounting. Using the PyCaret library (Ali, 2020), we automated model selection and determined that the random forest was the most effective algorithm for classifying vergence gestures.
@@ 4.1. Experimental setup
This study was approved by our Institutional Review Board. We recruited 20 participants (aged 25.75 ± 11.1), including 4 with normal vision and 16 with common refractive errors (12 myopia, 2 hyperopia, and 2 presbyopia). No participants reported other ocular-related disorders or convergence insufficiency.
@@
All the experiments were conducted in a meeting room (Fig. 10). To trigger the six vergence gestures, we used a Brock String (Hassan et al., 2017), a clinical tool commonly used to assess binocular vision and diagnose convergence insufficiency. Our string includes three colored beads positioned at the designated vergence distances: red at 30 cm, green at 70 cm, and yellow at 200 cm.
@@ 4.2. Data collection protocol
Following the Brock String exercise procedure (Scheiman and Wick, 2008), participants were instructed to perform vergence gestures such that the target bead remained in sharp focus while the non-target beads appeared blurred. Our experimental protocol (Fig. 9) consisted of two sessions, each comprising ten rounds. In each round, participants performed all six vergence gestures once, cued by an audible beep, with a three-second window to complete each movement. After ten rounds, participants removed the glasses for a ten-minute break before remounting the device. For each participant, we disinfected all recording apparatus in accordance with institutional hygiene protocols.
@@ 5.1. Vergence classification
Within-Session Performance. We performed a within-session evaluation for each of the 20 participants in our study (Fig. 11a, Table 3), and achieve an average accuracy of 82.68 ± 4.29% across all participants at classifying between the 6 different vergence gestures. We perform a per-class analysis of these results as a confusion matrix (Fig. 12(a)). The matrix shows that near-to-far and far-to-near vergence movements are never confused, which is due to their distinctively different positive-going and negative-going shapes with opposing polarity. There is virtually no misclassification between the 30 $\rightarrow$ 200 cm and 70 $\rightarrow$ 200 cm transitions, or between the 200 $\rightarrow$ 70 cm and 200 $\rightarrow$ 30 cm transitions, which is partly due to the wider separation between their EOG curves. However, the 30 $\rightarrow$ 70 cm movement creates an EOG waveform that can overlap with the 30 $\rightarrow$ 200 cm and 70 $\rightarrow$ 200 cm movements, and the 70 $\rightarrow$ 30 cm movement likewise can overlap between the 200 $\rightarrow$ 70 cm and 200 $\rightarrow$ 30 cm movements.
We also consider an alternate version of our system using only four gestures which have a high degree of separability: (30 $\rightarrow$ 200 cm, 70 $\rightarrow$ 200 cm, 200 $\rightarrow$ 30 cm, and 200 $\rightarrow$ 70 cm). Our evaluation (Fig. 13a, Table 3) shows that the system achieves a within-session classification accuracy of 97.99%, and achieves near-perfect discrimination as shown in the confusion matrix in Fig. 12c.
Figure 11. (a) Classification accuracy ($n=20$) for within-session (training and testing on data collected without remounting the glasses) versus cross-session (training on one session and testing on the unseen session after re-mounting glasses). P19 and P20 withdrew before session 2 due to time constraints. (b) Within-session 5-fold cross-validation accuracy per participant ($n=9$) for 2-, 3- and 4-channel configurations.
Cross-Session Performance. When the user remounts the glasses, slight shifts in electrode placement introduce signal drift. To evaluate the robustness of our system against remounting, we trained a model on one of the user’s data collection sessions and tested it on the other unseen session. We then reversed the roles and averaged the results. Our evaluation (Fig.  11b, Table 3) shows that system performance reduces modestly from 82.68% to 77.78% after remounting. The confusion matrix in Fig. 12(b) shows increased misclassification between same-direction vergence movements. This is because remounting induces data drift. However, for the 4-gesture version of our system, cross-session performance remained stable at 97.78%, suggesting that signal drift from remounting was not substantial enough to impact classification accuracy.
Cross-User Performance. Here, we evaluate the performance of a pretrained model in a leave-one-user-out setup across the 11 participants in our study, where no additional user enrollment and calibration is done. Our results in Table 3 show that in this scenario, our system achieves an accuracy of 77.36% and 97.43% without any new-user enrollment for the 6-class and 4-class version of our system respectively. Additionally, we show in Fig. 13a that these accuracies surpass the personalized model trained on 3.5 minutes of new-user data, which suggests that our system can be rapidly deployed to new users without a significant decrease in system performance.
@@
6-gesture accuracy 82.68 $\pm$ 4.29% 77.78 $\pm$ 6.17% 77.36 $\pm$ 6.17%
4-gesture accuracy 97.99 $\pm$ 1.96% 97.78 $\pm$ 2.02% 97.43 $\pm$ 2.06%
@@
Effect of Vertical Channels. We assessed the effect of vertically placed electrodes on system performance across 9 participants. Fig. 11b shows the per-participant accuracies for within-session evaluation for our proposed two‐channel configuration, a three‐ and four-channel configuration with an additional electrode pair above and below the left and right eye. 5-fold cross-validation shows that our two-channel configuration achieved 83.51% accuracy across six gestures, while the three- and four-channel configurations result in modest performance gains of 3.06% and 3.80% and accuracies of 86.57% and 87.31% respectively. These modest gains are likely because vergence is primarily, though not exclusively, a horizontal eye movement. However, all participants reported discomfort with the 3- and 4-channel configurations during data collection. Given these persistent ergonomic concerns, we excluded these configurations from the protocol for the remaining 11 participants. These findings further justify our decision to prioritize a two-channel design for the final system.
@@ 5.2. Subgroup analysis
To evaluate the generalization and robustness of VergeIO across diverse user groups, we conducted a subgroup analysis on the within-session data across 2 sessions. We assessed classification performance for both 6-gesture and 4-gesture tasks across demographics and physiological conditions, as shown in Fig. 14.
Sex. When comparing the classification performance by sex, the female participants ($n=8$) achieved a mean accuracy of 81.56 $\pm$ 4.67% for 6-gesture task and 97.81 $\pm$ 2.29% for 4-gesture task. The male participants ($n=12$) achieved 83.54 $\pm$ 3.98% and 98.75 $\pm$ 1.77% for 6-gesture and 4-gesture tasks, respectively.
Vision conditions. Participants who normally wear corrective lenses were instructed to wear them with our VergeIO glasses on top of them during the study to ensure they could clearly focus on vergence targets at 30, 70 and 200 cm. The mean classification accuracies for the tasks of 6-class and 4-class, respectively, were as follows: 84.03 $\pm$ 4.92% and 98.96 $\pm$ 1.29% for myopia ($n=12$); 82.91 $\pm$ 1.98% and 99.38 $\pm$ 1.25% for normal vision ($n=4$); 79.58 $\pm$ 0.60% and 96.88 $\pm$ 2.65% for hyperopia ($n=2$); and 77.90 $\pm$ 1.77% and 94.38 $\pm$ 0.88% for presbyopia ($n=2$).
Figure 13. (a) Effect of training duration on classification accuracy. (b) Motion artifact removal performance across vergence, motion, and noise EOG segments. (c) False positive rate under different movement scenarios: static, walking, chewing, and talking. A preamble consisting of an eyebrow raise to activate the system reduces the false positives to 0 in all four conditions.
Figure 14. Subgroup analysis on within-session classification accuracy for (a) 6-gestures and (b) 4-gestures.
@@ 5.3. Baseline comparison
We perform a comparison against two baseline EOG sensing systems. First, we compare our system with Mishra et al.’s work (Mishra et al., 2020), which is focused on vergence sensing using custom aerosol-jet–printed, stretchable, skin-like biopotential electrodes using a 2, 3, and 4-channel configurations. Their 2-channel configuration differs from ours and has a horizontal channel across both eyes, a vertical channel above and below the right eye, and a ground electrode obtrusively placed on the forehead, and reports an accuracy of 34%. In contrast, our system uses an unobtrusive 2-channel configuration using standard Ag/AgCl electrodes with the ground on the mastoid and achieves an accuracy of 82.68%. While their 3-channel and 4-channel configurations have increased accuracies of 87 and 91% respectively, additional electrodes bring about discomfort which negatively affect the user experience.
Mishra et al. (Mishra et al., 2020) VergeIO (ours)
Number of electrodes 5 - 7 4
Number of channels 2 - 3 2
In-session error 34% - 91% 82.68%
Cross-session error Not specified 77.78%
Cross-user error Not specified 77.36%
@@
Table 4. Comparison between our system and the baseline proposed by Mishra et al.  (Mishra et al., 2020)
Second, we compare our system’s performance against the EOG electrode layout in the commercial JINS MEME glasses (jin, 2026) (Fig. 4). Here, one participant (P4) wore electrodes at the union of both our system’s and JINS MEME’s electrode positions. We then collected data across ten rounds, each including all six vergence gestures. Five‐fold cross‐validation was performed independently on both setups. Our system achieves an accuracy of 88.33 $\pm$ 1.13%, while the JINS MEME achieves a lower accuracy of 50.0 $\pm$ 1.29%. This is in line with our prior SNR analysis showing that our vergence waveforms are on average 7.3 dB higher than those from JINS MEME (Fig. 5).
5.4. Compatibility with traditional EoG-based eye movements detection
While our primary goal of designing VergeIO is for the precise detection of vergence, we also evaluated the system’s compatibility with traditional EoG-based eye movements sensing. We collect a dataset including the four vergence gestures with high separability, along with seven additional eye movements: left and right winks, blinks, and four directional gazes (up, down, left, right). We collected 14 recordings for each class, and performed 5-fold cross validation. Our system achieved an accuracy of 96.1 $\pm$ 2.4%.
A few misclassifications were observed between upward and downward gaze, which is expected as our design intentionally omits vertical electrode channels to avoid user discomfort. However, our system still captures some vertical eye movement as we intentionally incorporate vertical separation between positive electrodes and reference electrode, as shown in Fig. 6b. These results show that our electrode configuration does not tradeoff accuracy in detecting eye movements sensed in prior EOG systems, and can detect them robustly along with vergence gestures, expanding detectable eye movements and possible applications on smart glasses as we will discuss in Sec. 6.